My Determine whether the given matrix is diagonalizable; if so, find a matrix P and a diagonal matrix D such that A - PDP1. (If the matrix is not dlagonalizable, enter DNE in any cell.) T o 1 0 A-1 20 L-1 1 1 [PD] Additional Materials Tutorial Show My Work (optiena) Submit Answer Save Progress Practice Another Version 25

linear algebra
Answer #1

Solution :

0 Diagonalize-1 2 0

\mathrm{Steps\:}:

\begin{pmatrix}0&1&0\\ -1&2&0\\ -1&1&1\end{pmatrix}

\mathrm{A\:matrix\:}A\mathrm{\:can\:be\:diagonalized\:if\:there\:exists\:an\:invertible\:matrix\:}P\mathrm{\:and\:diagonal\:matrix\:}D

\mathrm{\:such\:that\:}A=PDP^{-1}

\mathrm{Eigenvalues\:for\:}\begin{pmatrix}0&1&0\\ -1&2&0\\ -1&1&1\end{pmatrix}:\quad

\begin{pmatrix}0&1&0\\ -1&2&0\\ -1&1&1\end{pmatrix}

\mathrm{Eigenvalues\:of\:}A\mathrm{\:are\:the\:roots\:of\:the\:characteristic\:equ.\:}\det \left(A-\lambda \:I\right)=0

\det \left(\begin{pmatrix}0&1&0\\ -1&2&0\\ -1&1&1\end{pmatrix}-\lambda \begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix}\right)

=\det \left(\begin{pmatrix}0&1&0\\ -1&2&0\\ -1&1&1\end{pmatrix}-\begin{pmatrix}\lambda &0&0\\ 0&\lambda&0\\ 0&0&\lambda\end{pmatrix}\right)

=\det \begin{pmatrix}-\lambda &1&0\\ -1&2-\lambda&0\\ -1&1&1-\lambda\end{pmatrix}

\det \begin{pmatrix}-\lambda &1&0\\ -1&2-\lambda&0\\ -1&1&1-\lambda\end{pmatrix}

=-\lambda* \det \begin{pmatrix}2-\lambda&0\\ 1&1-\lambda\end{pmatrix}-1*\det \begin{pmatrix}-1&0\\ -1&1-\lambda \end{pmatrix}+0*\det \begin{pmatrix}-1&2-\lambda \\ -1&1\end{pmatrix}

=-\lambda* (\left(2-\lambda\right)\left(1-\lambda\right)-0*1)-1*(\left(-1\right)\left(1-\lambda\right)-0* \left(-1\right))+0*(\left(-1\right)*1-\left(2-\lambda\right)\left(-1\right))

=-\lambda* (\left(-\lambda+2\right)\left(-\lambda+1\right))-1*(-\left(-\lambda+1\right))-0* \left(-1\right))+0*(-1+2-\lambda)

=-\lambda* (\lambda^{2}-1\lambda-2\lambda+2*1)-1*(\lambda-1)+0*(-1+2-\lambda)

=-\lambda* (\lambda^2-3\lambda+2)-1*(\lambda-1)

=-\lambda ^3+3\lambda^2-2\lambda-\lambda+1

=-\lambda ^3+3\lambda^2-3\lambda+1

\mathrm{Solve\:}\:-\lambda ^3+3\lambda ^2-3\lambda +1=0:

-\lambda ^3+3\lambda ^2-3\lambda +1=0

-(\lambda ^3-3\lambda^2+3\lambda-1)=0

\mathrm{Apply\:cube\:of\:difference\:rule:\:}a^3-3a^2b+3ab^2-b^3=\left(a-b\right)^3

-(\left(\lambda -1\right)\left(\lambda-1\right)\left(\lambda-1\right))=0

-\left(\lambda -1\right)^3=0

\mathrm{Using\:the\:Zero\:Factor\:Principle:}

\mathrm{Solve\:}\:\lambda -1=0:

\lambda -1=0

\lambda =1

\mathrm{The\:final\:solution\:to\:the\:equation\:is:}

\lambda =1

\mathrm{The\:eigenvalues\:are:}

=1

\mathrm{Now,\:the\:eigenvectors\:for\:}\begin{pmatrix}0&1&0\\ -1&2&0\\ -1&1&1\end{pmatrix}:

\begin{pmatrix}0&1&0\\ -1&2&0\\ -1&1&1\end{pmatrix}

\mathrm{Solve\:}\:\left(A-\lambda \:I\right):\:\begin{pmatrix}0&1&0\\ -1&2&0\\ -1&1&1\end{pmatrix}-1\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix}

\begin{pmatrix}0&1&0\\ -1&2&0\\ -1&1&1\end{pmatrix}-1* \begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix}

=\begin{pmatrix}0&1&0\\ -1&2&0\\ -1&1&1\end{pmatrix}-\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix}

=\begin{pmatrix}0-1&1-0&0-0\\ \left(-1\right)-0&2-1&0-0\\ \left(-1\right)-0&1-0&1-1\end{pmatrix}

\mathrm{Simplify\:each\:element}

=\begin{pmatrix}-1&1&0\\ -1&1&0\\ -1&1&0\end{pmatrix}

\mathrm{Reduce\:}\begin{pmatrix}-1&1&0\\ -1&1&0\\ -1&1&0\end{pmatrix}:\quad

\begin{pmatrix}-1&1&0\\ -1&1&0\\ -1&1&0\end{pmatrix}

\mathrm{Cancel\:leading\:coefficient\:in\:row\:}\:R_2\:\mathrm{\:by\:performing}\:R_2\:\leftarrow \:R_2-1*R_1

=\begin{pmatrix}-1&1&0\\ 0&0&0\\ -1&1&0\end{pmatrix}

\mathrm{Cancel\:leading\:coefficient\:in\:row\:}\:R_3\:\mathrm{\:by\:performing}\:R_3\:\leftarrow \:R_3-1*R_1

=\begin{pmatrix}-1&1&0\\ 0&0&0\\ 0&0&0\end{pmatrix}

\mathrm{Multiply\:matrix\:row\:by\:constant:}\:R_1\:\leftarrow \:-1*R_1

=\begin{pmatrix}1&-1&0\\ 0&0&0\\ 0&0&0\end{pmatrix}

\mathrm{The\:system\:associated\:with\:the\:eigenvalue\:}\lambda =1

0 0 0/ z

\mathrm{This\:reduces\:to\:the\:equation}

x-y=0

\mathrm{Note\:that\:}z\mathrm{\:did\:not\:show\:up\:in\:the\:equations\:which\:means\:}\begin{pmatrix}1&-1&0\\ 0&0&0\\ 0&0&0\end{pmatrix}\begin{pmatrix}0\\ 0\\ z\end{pmatrix}=\begin{pmatrix}0\\ 0\\ 0\end{pmatrix}\mathrm{\:regardless\:of\:the\:value\:of\:}z

\mathrm{Therefore,\:the\:eigenvectors\:associated\:with\:the\:eigenvalue\:}\lambda =1

\begin{pmatrix}0\\ 0\\ 1\end{pmatrix}

\mathrm{Isolate}

x=y

\mathrm{Plug\:into\:}\begin{pmatrix}x\\ y\\ z\end{pmatrix}

v=\begin{pmatrix}y\\ y\\ 0\end{pmatrix}\space\space\:y\ne \:0

\mathrm{Let\:}y=1

\begin{pmatrix}1\\ 1\\ 0\end{pmatrix}

\mathrm{All\:eigenvectors\:for\:the\:eigenvalue\:}\lambda =1

\begin{pmatrix}0\\ 0\\ 1\end{pmatrix},\:\begin{pmatrix}1\\ 1\\ 0\end{pmatrix}

\mathrm{To\:allow\:diagonalization,\:the\:number\:of\:Eigenvectors\:must\:equal\:the\:matrix\:dimensions.}

\mathrm{There\:are\:}2\mathrm{\:Eigenvectors\:which\:is\:less\:then\:}3\mathrm{\:and\:therefore\:the\:matix\:cannot\:be\:diagonalized}

=\mathrm{Not\:diagonizable}